Solve the given integral equation or integro-differential equation for y(t). t y(t)+27 [ct –v)y(v) dv= 4t, y(0)=0 0 y(t) =
Answer #1

22 y(0) 20 4647 4 22.5-vy d = 4,5 400) Leibnitz Integral Rule y! (0)+27x0 20 als fxstydt) = f(x, 6(x)). I b (x) - f (x, a (x2) ₂20 y(t) = (ket + GEL), -32 2) y (t) = ((2 t + g +2) e 3+ (-3) test (a + 2tc,) 2) y(0), (0+o) (-3) e 3x0 + e 30 (+0) ₃0 2